A machine fills coffee powder in pouches, with an average of 200 gm and a standard deviation of 4 gm. Assuming that the coffee weight is normally distributed. Find the probability that a coffee pouch selected at random will contain the following quantity of a coffee: I) At least 200 gm. II) Between 200 to 206 gm


1
Expert's answer
2022-04-06T09:38:14-0400

mean("\bar x" )=200

standard deviation (s)=4

z="\frac{x-\bar x}{s}"

1.

x=200

P(x"\geq200)=1-P(x<200)"

z="\frac{200-200}{4}" =0

P(z<0)=0.5

P(x"\geq" 200)=1-0.5

P(x"\geq" 200)=0.5

The probability that a coffee pouch selected at random will contain at least 200 gm is 0.5.

2. Between 200 and 206 .

P(200<x<206)= P(z1<x<z2)

z2="\frac{206-200}{4}"

z2=1.5

P(z2"\le" 1.5)=0.9332

using z1 previously computed

P(z1"\le"0)=0.5

P(z1<x<z2)=0.9332-0.5

P(z1<x<z2)=0.4332

The probability that a coffee pouch selected at random will contain between 200 gm and 206 gm is 0.4332.